Explain the meaning of the term “variable”. Write a puzzle for a 10-year child. Write an equation with a variable in it which gives a solution to the question raised by the puzzle. Also give a solution to the puzzle and interpret it.
1
Expert's answer
2020-03-30T10:00:37-0400

Variable is a designation of an unknown object in the task to be found.

Variables are very common in our life, but we rarely realize it.

For example:

Martin wants to buy a lollipop. It costs 5$. But Martin have only 2$. How much dollars he needs to buy a lollipop.


Solution:

We know that Martin has 2$. But the prise is 5$. Let's write variables like "x". We have to find, how much he needs.


2 (how much he has now) + x(how much he needs) = 5.

"2 + x = 5"

Let's subtract 2 from the left and from the right

"x = 5 - 2"

--> "x = 3"

Martin needs extra $3 to buy a lollipop.